OpenStudy (anonymous):

[6.02] Use the Substitution Method to solve the following system of equations. 2x - 9y = 1 x - 4y = 1 (1, 5) (-5, -1) (5, 1) (-5, 1)

hartnn (hartnn):

do you know substitution method ? and tried this ?

OpenStudy (jack1):

hey jenn, how're ya doin? so you have 2 equations, right?

OpenStudy (anonymous):

yeah

OpenStudy (jack1):

2x - 9y = 1 EQN 1 x - 4y = 1 ENQ 2 and the substitution method means you rearrange one of the equations to get it in the from of 1 variable (ie x ) = number +/- other variable (ie y) so rearranging equation 2 x - 4y = 1 ENQ 2 (so add 4y to both sides to get x by itself) x - 4y +4y = 1 + 4y x = 1 + 4y new EQN 2 or EQN 2a if you like

OpenStudy (anonymous):

So whats the answer?

OpenStudy (jack1):

now equation 2a explains what x is so now you can substitute that back into equation 1 2x - 9y = 1 and you know x = 1 + 4y so 2(1 + 4y) - 9y = 1 2 + 8y - 9y = 1 can you solve for y from here...?

OpenStudy (anonymous):

this is so confusing -.-

OpenStudy (anonymous):

Add together the common variables

OpenStudy (anonymous):

-y?

OpenStudy (anonymous):

y= -1 ?

OpenStudy (jack1):

2 + 8y - 9y = 1 minus 2 from both sides 2 -2 + 8y - 9y = 1-2 -y = -1 so y = 1

OpenStudy (anonymous):

i get it now, thanks!

OpenStudy (jack1):

cool, so u can solve for x now?

OpenStudy (anonymous):

5

OpenStudy (jack1):

spot on dude!